One way the U.S. Environmental Protection Agency (EPA) tests for chloride contaminants in water is by titrating a sample of silver nitrate solution. Any chloride anions in solution will combine with the silver cations to produce bright white silver chloride precipitate. Suppose an EPA chemist tests a 250. mL sample of groundwater known to be contaminated with tin(II) chloride, which would react with silver nitrate solution like this:

SnCl 2 (aq) + 2AgNO3 (aq) → 2 AgCl (s) + Sn(NO3)2 (aq)

The chemist adds 56.0 mM silver nitrate solution to the sample until silver chloride stops forming. He then washes, dries, and weighs the precipitate. He finds he has collected 5.2 mg of silver chloride.

Calculate the concentration of tin(II) chloride contaminant in the original groundwater sample. Be sure your answer has the correct number of significant digits.
Answer #1

Moles of AgCl = mass of salt / molar mass = 5.2 * 10-3 / 143.32 = 3.63 * 10-5 moles

From the equation, moles of Sn = moles of AgCl / 2 = 3.63 * 10-5 / 2 = 1.82 * 10-5 moles

Concentration = moles / Volume in liters = 1.82 * 10-5 / 0.25 = 7.26 * 10-5 M